# ByteCheckpoint: A Unified Checkpointing System for Large Foundation Model Development

### v2

@@ Abstract
Checkpointing to preserve training states is crucial during the development of Large Foundation Models (LFMs), for training resumption upon various failures or changes in GPU resources and parallelism configurations. In addition, saved checkpoints are dispatched to evaluation tasks or transferred across different training stages (e.g., from pre-training to post-training). All these scenarios require resharding distributed checkpoints from one parallelism to another. In production, different LFMs are trained with various frameworks and storage backends, depending on model sizes and training scales. A high-performance checkpointing system is needed to enable efficient checkpoint management at scale. This paper presents ByteCheckpoint, an industrial-grade checkpointing system for large-scale LFM training. ByteCheckpoint employs a parallelism-agnostic checkpoint representation that enables efficient load-time checkpoint resharding. ByteCheckpoint advocates a generic checkpoint saving/loading workflow to accommodate multiple training frameworks and support different storage backends. To ensure high I/O efficiency, we take a full-stack approach to optimize saving/loading plan generation, critical stages of checkpointing pipelines, and irregular tensor processing required by resharding. To guarantee the scalability of ByteCheckpoint in large-scale training, we enhance the storage system to efficiently handle high volumes of checkpointing I/O requests, devise communication optimizations within the checkpointing workflow, and introduce a suite of monitoring tools to analyze performance and detect bottlenecks. Compared to existing open-source checkpointing systems [46, 40], ByteCheckpoint significantly reduces runtime checkpoint stalls, achieving an average reduction of 54.20$\times$. For saving and loading times, ByteCheckpoint achieves improvements of up to 9.96$\times$ and 8.80$\times$, respectively.
@@ 1 Introduction
Large Foundation Models (LFMs) in language [5, 11, 34], vision [35, 31], audio [47] and other modalities are revolutionizing today’s AI landscape, spawning a variety of downstream applications such as conversational agents [44, 38], coding assistants [39], painting tools [41], and video [43] generators.
Unlike traditional deep learning model training, modern LFM training is significantly more complex. It has multiple training stages, including pre-training [20] and post-training [11, 45]. Furthermore, evaluation tasks are integrated into these stages to effectively assess model quality. Beyond the complexity, the development of LFMs is notably resource-intensive and time-consuming, largely due to immense model sizes (e.g., the largest version of Llama 3.1 comprises 405 billion parameters) and massive training datasets. The scale of LFM training can even be up to running on 12,288 GPUs [20].
@@
First, efficient and unified checkpoint management is necessary throughout the life cycle of real-world LFM development. During training, checkpoint resharding is commonly required. This procedure transforms saved distributed checkpoints so they can be correctly loaded into a new parallelism configuration that differs from the one used for their creation. In pre-training, variations in parallelism occur when problematic machines are removed, the available GPU quota fluctuates [56, 23], or training configurations (e.g., context length) and system optimization techniques [20] (e.g., kernel fusion, computation and communication overlapping ) are adjusted. Across different stages and tasks, parallelism varies according to the scales of resources and datasets in use. Highly efficient checkpoint resharding is needed to minimize extra overhead and maximize the end-to-end Effective Training Time Ratio (ETTR, calculated as the ratio between the actual training time and the total job running time). Moreover, on an industrial AI platform, LFM training jobs are initiated by various internal users or cloud customers who may choose different training frameworks (e.g., Megatron-LM [36], PyTorch FSDP [57], DDP [24], and others [50]) and select storage backends such as local disk, Hadoop Distributed File System (HDFS), or Network-Attached Storage (NAS) for storing checkpoints according to job characteristics and personal preferences. Crafting customized checkpointing modules and workflows to accommodate these ad-hoc implementations complicates system development and substantially increases maintenance costs. It is vital to provide generic workflows for different training frameworks and storage backends.
@@
Most of the existing checkpointing systems [27, 12, 52, 15] assume consistent parallelism during the saving and loading of checkpoints, failing to address the demands for checkpoint resharding. Although some efforts from the open-source community are devoted to developing checkpointing systems capable of resharding, they have several limitations. Some perform resharding in an inefficient offline manner [25], while others only support specific training frameworks and storage backends [46, 40]. Moreover, without customized optimizations and deployment experience, they suffer from sub-optimal I/O performance and lack the scalability to support large-scale LFM training in real-world production.
This paper presents the design, implementation, and engineering experience of ByteCheckpoint, a checkpointing system crafted for LFM development. ByteCheckpoint incorporates a unified architecture (Sec. 3.1) featuring a resharding-on-loading mechanism for distributed checkpoints (load-time checkpoint resharding) and supports multiple training frameworks and storage backends. Besides, it integrates full-stack I/O performance and scalability optimization techniques.
$\triangleright$ ByteCheckpoint’s checkpoint storage representation is decoupled from the specific parallelism adopted during training (Sec. 3.2), enabling efficient load-time checkpoint resharding. For model and optimizer state representation, we separate each tensor shard’s metadata from its numerical values and consolidate all the metadata into one global file. The metadata of a tensor shard includes its basic runtime, position, and storage information. We unify the position information of the tensor shard into the index tuple (fqn, nD_offsets, nD_lengths). For dataloader state representation, we divide dataloader states into replicated and sharded states, storing unique sharded states in individual files while saving the replicated one only by the training worker whose global rank is 0. File index information of sharded states is also recorded in the global metadata file.
@@
$\triangleright$ ByteCheckpoint implements multiple optimizations (Sec. 4) to enhance I/O performance, including balanced and zero-redundancy plan generation, fully asynchronous execution pipelines, and efficient irregular tensor processing. We share our engineering experience in optimizing storage systems to support massive I/O requests, optimizing collective communications to guarantee stability, and designing monitoring and visualization tools for performance analysis (Sec. 5). This full-stack approach scales ByteCheckpoint to support the training of a 405B LFM on 8,960 GPUs while still maintaining high efficiency.
ByteCheckpoint is deployed on our industrial AI platform with tens of thousands of GPUs for various LFM training tasks, including pre-training and post-training of language models, multi-modal understanding and generation models (e.g., for video generation). ByteCheckpoint demonstrates significant advantages over existing open-source checkpointing systems, including PyTorch DCP [46] and Megatron Distributed Checkpoint [40] (MCP). Compared to the baselines, ByteCheckpoint achieves improvements ranging from 12.13$\times$ to 161.50$\times$ in terms of checkpoint stalls reduction, making the end-to-end checkpoint saving and loading procedures 6.05$\times$ and 3.88$\times$ faster on average, respectively.
@@ 2.1 LFM development
Production pipeline. As depicted in Fig. 1, the development of LFM comprises pre-training and post-training stages [11]. During the initial pre-training phase, The LFM is iteratively trained on extensive data collected from multiple sources to absorb knowledge about the world. Subsequently, continual pre-training is employed to enhance the foundation model’s capabilities. For instance, the pre-training of large language models (LLMs) typically involves long-context continual training to gradually increase the supported context length of LLMs while still maintaining their short-context ability. Post-training is employed to align the pre-trained model with human feedback or generate content that adheres to given instructions [54]. Various task-specific labeled datasets (e.g., multilingual, code, math, reasoning, etc.) are involved in fine-tuning the LFMs, followed by reinforcement learning, which includes the reward modeling and then performing Proximal Policy Optimization (PPO), or directly conducting Direct Preference Optimization [32] (DPO). Due to the reduced size of these datasets, relatively fewer GPUs are involved in post-training. During these stages, automatic evaluation [18, 12] is periodically triggered to get intermediate model checkpoints and assess model quality with diverse criteria.
@@
Figure 1: An overview of the LFM training pipeline.
@@ 2.2 LFM Checkpointing Requirements
Load-time checkpoint resharding. During the life cycle of LFM development, checkpoint resharding is consistently required due to changes in training parallelism under different scenarios (Fig. 2). For pre-training, the allocated GPU quota can vary due to removal of faulty machines [20, 18], adding new machines released from completed tasks, or running on tidal resources [56, 23]. It is often necessary to adjust training parallelism in response to the resource changes, to maximize resource utilization. For example, in Fig. 2, eight distributed checkpoint files are initially stored, then loaded into six training workers upon resuming. Additionally, continual pre-training such as long-context training changes the context length, which also requires GPU quota or parallelism adjustment. Moreover, at the onset of large-scale pre-training, AI engineers often need to experiment with various model configurations and optimization techniques [20], since conclusions from small-scale profiling or simulations do not always translate to optimal performance in large-scale training. In this phase, adjusting parallelism configurations is common, resulting in frequent training resumption. When transitioning from pre-training to post-training , the number of GPUs involved typically decreases due to reduced training data in the latter. Distributed checkpoints saved from pre-training are frequently resharded to align with the specific workloads of each post-training task. As depicted in Fig. 2, only 4 GPUs are involved for a fine-tuning task in the post-training stage, so the checkpoints are resharded accordingly. Evaluation tasks in both stages require loading model checkpoints and conducting model inference on separate GPU resources; their parallelism needs to be adjusted to align with assigned GPUs used for specific evaluation datasets. Fig. 2 shows 4-GPU eval task requesting model states from pre-training checkpoints.
We collected the number of times of checkpoint resharding demands on our AI platform (training text, image, and video generation models ) over the past six months and identified 1,870 instances of checkpoint resharding during pre-training resumption, 19,844 for running evaluation tasks, and 13,080 due to cross-stage reconfiguration.
As a critical step in the life cycle of LFM development, the efficiency of checkpoint resharding is vital to minimize extra overheads. In our AI platform, the previous common practice is to develop offline checkpoint resharding scripts and adapt the scripts whenever a new checkpoint resharding scenario arises. Examples of offline resharding with customized scripts are given in Fig. 3. Customizing offline scripts is labor-intensive. For GPU states resharding, the offline scripts must cover all different components in the LFM and its optimizer, and cater to the diverse behavior of each component under different parallelism strategies. For example, with tensor parallelism (TP), GEMM operators in attention and MLP blocks are sharded along different dimensions, while other operators like LayerNorm [7] are replicated across GPUs. When hybrid 3D parallelism [28] is adopted, TP-sharded tensors of one layer (module) in the distributed optimizer are first flattened and then merged before being sharded according to the designated data parallelism (DP) degree. Offline scripts must implement resharding logic that is tightly coupled with combinations of model (optimizer) components and parallelism strategies. Additionally, special algorithmic optimization techniques such as GQA [19] and MLA [58] change the tensor layout of certain operators (e.g., the query-key-value projection GEMM operator in the attention block), necessitating corresponding resharding support. To handle various cases in our production environment, our largest script even includes 3193 lines of Python code. This complexity results in significant engineering efforts for both development and maintenance.
@@
In addition to the development costs, running resharding scripts leads to a substantial waste of GPU time and resources. Table 1 presents the cost of executing resharding jobs in various resharding scenarios. Before training can resume or new evaluation tasks can begin, independent jobs that execute the resharding scripts must be submitted in advance. These jobs download checkpoints from the storage systems, reshard distributed checkpoints to given parallelism configurations and upload new checkpoints back to the storage systems. The targeted training or evaluation jobs cannot be executed until the completion of resharding jobs, resulting in prolonged pending time. Moreover, since checkpoints created by resharding scripts are coupled with specific parallelism, they cannot be reused freely, which in turn increases the storage overhead.
@@
Training Resuming Evaluation Cross-stage Transition
1870.38s 593.21s 650.34s
@@
Multiple frameworks and storage backends. On our AI platform, a wide range of training frameworks are used, such as Megatron-LM [36], DDP [24], FSDP [57], veScale [20, 50], etc. Table 2 lists the top three preferred training frameworks on our platform, based on six months of trace analysis. Users typically adopt Megatron-LM [36] for training large language foundation models. FSDP [57] is preferred for training tasks involving text-to-video or text-to-speech models, and DDP [24] is commonly used to train image encoder components of multimodal foundation models or for routine algorithm testing. In addition, users can choose from various storage backends for checkpoint persistence depending on the scenario, ranging from local disks for debugging to HDFS or NAS for formal training tasks. Each training framework comes with its own checkpoint module, file format, and ad-hoc implementation of the save/load logic. Tailoring optimized checkpoint management to each framework would be labor-intensive, leading to fragmented checkpointing interfaces across training frameworks and storage backends, complicating the codebase and increasing maintenance costs. Therefore, it is vital to provide generic workflows for different training frameworks and storage backends.
@@
Figure 4: Checkpointing efficiency impacts failure recovery and evaluation tasks. D2H here denotes the Device-to-Host copy phase in checkpointing.
@@
Efficient and scalable I/O performance. Mainstream LFMs feature extensive model sizes, reaching hundreds of billions of parameters [11]. Consequently, the size of training states also increases significantly, imposing substantial overhead for checkpoint saving and loading. By analyzing our previous LFM training jobs, we observe that the average end-to-end time required to save checkpoints of a GPT 175B model, trained on 4096 GPUs, to HDFS can be 200 seconds. This duration substantially exceeds the time required for a single training iteration. Even though this time-consuming procedure can be partially removed from the critical path of model training by adopting asynchronous checkpointing [27, 20, 51, 26], it is still crucial to expedite checkpoint saving and loading, given that frequent failures are inevitable in large-scale training [20] (e.g., 466 job interruptions were observed during 54 days of training llama3.1 [11]). As illustrated in Fig. 4, though the checkpointing procedure overlaps with training, faster checkpointing allows more intermediate checkpoints to be stored before failure occurs, enabling training resumption from more recent states. In addition, evaluation tasks are triggered during training, and intermediate checkpoints are periodically pulled from the persistent storage for those tasks. Faster checkpointing allows timely execution of those tasks, reducing blocking time due to preparing the checkpoints they need.
@@
Scaling the checkpointing system while maintaining the high I/O performance is non-trivial. Sub-optimal and risky designs, which are hard to detect in small-scale settings, can lead to severe performance bottlenecks or even catastrophic job failures in large-scale training. For example, massive read/write requests for checkpoint files from the training cluster to the storage systems (e.g., HDFS) can overload the NameNode, causing delays in file metadata operations. Additionally, naive implementations of collective communications, such as integrity-checking barriers, introduce substantial initialization and synchronization overheads. These overheads can even result in communication timeouts, ultimately causing the entire training job to fail. Furthermore, as the number of GPUs scales up, analyzing system performance and detecting errors among training and I/O workers across multiple machines becomes increasingly challenging.
@@
Existing checkpointing systems, such as CheckFreq [27], Check-N-Run [12], and Gemini [52], operate under the assumption of consistent parallelism and do not address the need for checkpoint resharding. DCP [46] and MCP [46] incorporate checkpoint resharding capabilities but are limited in terms of supported parallelism strategies and training frameworks. Additionally, both systems exhibit suboptimal I/O performance and scalability issues. A comprehensive discussion with related works is provided in Sec 7. In ByteCheckpoint, we design decoupled storage representation for efficient load-time checkpoint resharding, propose generic saving/loading workflows to support multiple frameworks and storage backends, integrate full-stack optimizations to enhance I/O performance and share our experience in scaling the checkpointing system to support real-world LFM training.

### v3

@@ Abstract
Checkpointing to preserve training states is crucial during the development of Large Foundation Models (LFMs), for training resumption upon various failures or changes in GPU resources and parallelism configurations. In addition, saved checkpoints are dispatched to evaluation tasks or transferred across different training stages (e.g., from pre-training to post-training). All these scenarios require resharding distributed checkpoints from one parallelism to another. In production environments, different LFMs are trained with various frameworks and storage backends, depending on model sizes and training scales. A high-performance checkpointing system is needed to enable efficient checkpoint management at scale throughout the lifecycle of LFM development.
We introduce ByteCheckpoint, an industrial-grade checkpointing system for large-scale LFM training. ByteCheckpoint features: a parallelism-agnostic checkpoint representation that enables efficient load-time checkpoint resharding; a generic checkpoint saving/loading workflow to accommodate multiple training frameworks and support different storage backends; full-stack optimizations to ensure high I/O efficiency and scalability; a suite of monitoring tools to streamline large-scale performance analysis and bottleneck detection. Compared to existing open-source checkpointing systems [58, 52], ByteCheckpoint significantly reduces runtime checkpoint stalls, achieving an average reduction of 54.20$\times$. For saving and loading times, ByteCheckpoint achieves improvements of up to 9.96$\times$ and 8.80$\times$, respectively.
@@ 1 Introduction
Large Foundation Models (LFMs) in language [5, 13, 43, 33], vision [44, 40], audio [59] and other modalities are revolutionizing today’s AI landscape, spawning a variety of downstream applications such as conversational agents [56, 49], coding assistants [51], painting tools [53], and video [55] generators.
Unlike traditional deep learning model training, LFM training is significantly more complex. It has multiple training stages, including pre-training [9, 25] and post-training [67, 38, 46, 70]. Furthermore, evaluation tasks are integrated into these stages to effectively assess model quality. Beyond the complexity, the development of LFMs is notably resource-intensive and time-consuming, largely due to immense model sizes and massive training datasets (e.g., DeepSeek-V3 comprises 671 billion total parameters, and is pre-trained on 14.8 trillion tokens) [13, 33]. The scale of LFM training can even be up to running on 12,288 GPUs [25].
@@
First, efficient and unified checkpoint management is necessary throughout the life cycle of real-world LFM development. During training, checkpoint resharding is commonly required. This procedure transforms saved distributed checkpoints so they can be correctly loaded into a new parallelism configuration that differs from the one used for their creation. In pre-training, variations in parallelism occur when problematic machines are removed, the available GPU quota fluctuates [72, 29], or training configurations (e.g., context length) and system optimization techniques [25] (e.g., kernel fusion, computation and communication overlapping ) are adjusted. Across different stages and tasks, parallelism varies according to the scales of resources and datasets in use. Highly efficient checkpoint resharding is needed to minimize extra overhead and maximize the end-to-end Effective Training Time Ratio (ETTR, calculated as the ratio between the productive training time and the wallclock time of a job) [27]. Moreover, on an industrial AI platform, LFM training jobs are initiated by various internal users or cloud customers who may choose different training frameworks (e.g., Megatron-LM [47], PyTorch FSDP [73], DDP [31], and others [63]) and select storage backends such as local disk, Hadoop Distributed File System (HDFS), or Network-Attached Storage (NAS) for storing checkpoints according to job characteristics and personal preferences. Crafting customized checkpointing modules and workflows to accommodate these ad-hoc implementations complicates system development and substantially increases maintenance costs. It is vital to provide generic workflows for different training frameworks and storage backends.
@@
Existing checkpointing systems [35, 14, 66, 19] assume consistent parallelism, and fail to address the demands for checkpoint resharding. Although some efforts from the open-source community are devoted to developing checkpointing systems capable of resharding, they have several limitations. Some perform resharding in an inefficient offline manner [32], while others only support specific training frameworks and storage backends [58, 52]. Moreover, without customized optimizations and deployment experience, they suffer from sub-optimal I/O performance and lack the scalability to support large-scale LFM training in real-world production.
This paper presents the design, implementation, and deployment experience of ByteCheckpoint, a checkpointing system crafted for LFM development. ByteCheckpoint incorporates a unified architecture (Sec. 3.1) featuring an automatic resharding-on-loading mechanism for distributed checkpoints (load-time checkpoint resharding) and supports multiple training frameworks and storage backends. It integrates full-stack I/O performance and scalability optimization techniques.
$\triangleright$ ByteCheckpoint’s checkpoint representation is decoupled from the specific parallelism adopted during training (Sec. 3.2), enabling efficient load-time checkpoint resharding. For model and optimizer state representation, we separate each tensor shard’s metadata from its numerical values and consolidate all the metadata into one global file. The metadata of a tensor shard includes its basic runtime, position, and storage information. For the representation of dataloader states, we divide them into replicated and sharded states, storing sharded states in individual files while the replicated ones are only saved by the training worker whose global rank is 0.
@@
$\triangleright$ ByteCheckpoint implements multiple optimizations (Sec. 4) to enhance I/O performance, including balanced and zero-redundancy plan generation, fully asynchronous execution pipelines, and efficient irregular tensor processing. We share our engineering experience in optimizing storage systems to support massive I/O requests, optimizing collective communications to guarantee stability, and designing monitoring and visualization tools for performance analysis and bottleneck detection (Sec. 5). This full-stack approach scales ByteCheckpoint to support the training of a 405B LFM on 8,960 GPUs while still maintaining high efficiency.
ByteCheckpoint is deployed on our industrial AI platform with tens of thousands of GPUs for various LFM training tasks, including pre-training and post-training of language models, multi-modal understanding, and generation models (e.g., for video generation). ByteCheckpoint demonstrates significant advantages over existing open-source checkpointing systems, including PyTorch DCP [58] and Megatron Distributed Checkpoint [52] (MCP). Compared to the baselines, ByteCheckpoint achieves improvements ranging from 12.13$\times$ to 161.50$\times$ in terms of checkpoint stalls reduction, making the end-to-end checkpoint saving and loading procedures 6.05$\times$ and 3.88$\times$ faster on average, respectively.
@@ 2.1 LFM development
Figure 1: An overview of the training pipeline of LFM.
Production pipeline. As depicted in Fig. 1, the development of LFM comprises pre-training and post-training stages. During the initial pre-training phase, The LFM is iteratively trained on extensive data collected from multiple sources to absorb knowledge about the world. Subsequently, continual pre-training is employed to enhance the foundation model’s capabilities. For instance, the pre-training of large language models (LLMs) typically involves long-context continual training to gradually increase the supported context length of LLMs. Post-training is employed to align the pre-trained model with human feedback or enhance the model’s reasoning capabilities [57, 18]. Various task-specific labeled datasets (e.g., multilingual, code, math, reasoning, etc.) are involved in fine-tuning the LFMs, followed by reinforcement learning, typically including the reward modeling and then performing Proximal Policy Optimization [45] (PPO), or directly conducting Direct Preference Optimization [41] (DPO). Due to the reduced size of these datasets, relatively fewer GPUs are involved in post-training. Automatic evaluation [22, 14] is periodically triggered to get intermediate model checkpoints and assess the quality with diverse criteria.
@@
2.2 Checkpoint Resharding Scenarios
During the life cycle of LFM development, checkpoint resharding is consistently required due to changes in parallelism under different scenarios (Fig. 2):
(1) Training Resumption. GPU quota allocated for an LFM training job can vary due to the removal of faulty machines [25, 22], adding new machines released from completed tasks, or running on tidal resources [12, 72, 29, 68]. It is often necessary to adjust training parallelism in response to the resource changes, to maximize resource utilization. Additionally, at the onset of large-scale pre-training, AI engineers often need to experiment with various model configurations and optimization techniques [25], since conclusions from small-scale profiling or simulations do not always translate to optimal performance in large-scale training. In this phase, adjusting parallelism configurations is common, resulting in frequent training resumption. Moreover, long-context training changes the context length, which also requires GPU quota or parallelism adjustment. As shown in the example in Fig. 2, 8 distributed checkpoint files are initially stored, then loaded into 6 training workers upon resuming.
(2) Cross-Stage Transition. When entering post-training , the number of GPUs involved typically decreases due to reduced training data in the latter. Checkpoints saved from pre-training are frequently resharded to align with the specific workloads of each post-training task. As depicted in Fig. 2, only 4 GPUs are involved for a fine-tuning task in the post-training stage, so the checkpoints are resharded accordingly.
(3) Evaluation. Evaluation tasks in both stages require loading model checkpoints and conducting inference on separate resources; their parallelism needs to be adjusted to align with assigned GPUs used for specific datasets. Fig. 2 shows a 4-GPU eval task reshards model checkpoints from pre-training.
We collected the number of times of checkpoint resharding demands on our AI platform (training text, image, and video generation models) over the past six months and identified 1,870 instances of checkpoint resharding during pre-training resumption, 13,080 due to cross-stage reconfiguration, and 19,844 for running evaluation tasks.
@@ 2.3 LFM Checkpointing Requirements
Load-time checkpoint resharding. As a critical step in the life cycle of LFM development, the efficiency of checkpoint resharding is vital to minimize extra overheads. In our AI platform, the previous common practice is to develop offline checkpoint resharding scripts and adapt the scripts whenever a new checkpoint resharding scenario arises. This approach is inefficient and labor-intensive (see Appendix A for further details). In addition to the development costs, running resharding scripts leads to a substantial waste of GPU time and resources. Table 1 presents the cost of executing resharding jobs in various scenarios. Before training can resume or new evaluation tasks can begin, independent jobs that execute the resharding scripts must be submitted in advance. These jobs download checkpoints from the storage systems, reshard distributed checkpoints to given parallelism configurations and upload new checkpoints back to the storage systems. The targeted training or evaluation jobs cannot be executed until the completion of resharding jobs, resulting in prolonged pending time. Moreover, since checkpoints created by resharding scripts are coupled with specific parallelism, they cannot be reused freely, which in turn increases the storage overhead.
@@
Training Resumption Cross-Stage Transition Evaluation
1870.38s 650.34s 593.21s
@@
Multiple frameworks and storage backends. On our AI platform, a wide range of training frameworks are used, such as Megatron-LM [47], DDP [31], FSDP [73], veScale [25, 63], etc. Table 2 lists the top three preferred training frameworks on our platform, based on six months of trace analysis. Users typically adopt Megatron-LM [47] for training large language foundation models. FSDP [73] is preferred for training tasks involving text-to-video or text-to-speech models, and DDP [31] is commonly used to train image encoder components of multimodal foundation models or for routine algorithm testing. In addition, users can choose from various storage backends for checkpoint persistence depending on the scenario, ranging from local disks for debugging to HDFS or NAS for formal training tasks. Each training framework comes with its own checkpoint module, file format, and ad-hoc implementation of the save/load logic. However, these modules lack critical features for production, such as load-time resharding, asynchronous checkpointing, and support for remote persistent storage. Integrating these features for each framework’s checkpoint module and tailoring optimized implementations requires repeated engineering efforts. This leads to inconsistent checkpointing interfaces across training frameworks and storage backends, complicating the codebase. Moreover, the maintenance of diverse checkpoint file formats from different frameworks adds to the complexity of implementing checkpoint transfer logic across training stages and deploying models for evaluation and inference tasks. Therefore, it is vital to provide generic workflows for different training frameworks and storage backends.
@@
Figure 3: Checkpointing efficiency impacts failure recovery and evaluation tasks. D2H denotes the Device-to-Host copy.
@@
Efficient and scalable I/O performance. Mainstream LFMs feature extensive model sizes, reaching hundreds of billions of parameters [13]. Consequently, the size of training states also increases significantly, imposing substantial overhead for checkpoint saving and loading. By analyzing our previous LFM training jobs, we observe that the average end-to-end time required to save checkpoints of a GPT 175B model, trained on 4096 GPUs, to HDFS can be 200 seconds. This duration substantially exceeds the time required for a single training iteration. Even though this time-consuming procedure can be partially removed from the critical path of model training by adopting asynchronous checkpointing [35, 25, 65, 34], expediting end-to-end checkpointing remains crucial to minimize training progress loss [66] caused by inevitable frequent failures in large-scale training [25]. As depicted in Fig. 3, although checkpointing overlaps with training, its rapid completion allows more intermediate checkpoints to be stored before a failure occurs, enabling resumption from more recent states and improving ETTR. Besides, evaluation tasks are triggered during training, and intermediate checkpoints are periodically pulled for these tasks. Faster checkpointing ensures their timely execution, reducing blocking time due to preparing these checkpoints in remote persistent storage.
@@
Scaling the checkpointing system while maintaining the high I/O performance is non-trivial. Sub-optimal and risky designs, which are hard to detect in small-scale settings, can lead to severe performance bottlenecks or even catastrophic job failures in large-scale training. For example, massive read/write requests for checkpoint files from the training cluster to the storage systems (e.g., HDFS) can overload the master node, causing delays in file metadata operations. Additionally, naive implementations of collective communications, such as integrity-checking barriers, introduce substantial initialization and synchronization overheads. These overheads can even result in communication timeouts, ultimately causing the entire training job to fail. Furthermore, as training scales up, efficiently analyzing system performance and detecting errors among training and I/O workers across multiple machines becomes increasingly challenging.
@@
Existing checkpointing systems, such as CheckFreq [35], Check-N-Run [14], and Gemini [66], operate under the assumption of consistent parallelism and do not address the need for checkpoint resharding. DCP [58] and MCP [58] incorporate checkpoint resharding capabilities but are limited in terms of supported parallelism strategies and training frameworks, I/O performance, and scalability. A comprehensive discussion with related works is provided in Appendix F. In ByteCheckpoint, we design decoupled storage representation for efficient load-time checkpoint resharding, propose generic saving/loading workflows to support multiple frameworks and storage backends, integrate full-stack optimizations to enhance I/O performance and share our experience in scaling the checkpointing system to support real-world LFM training.